solve the system of equations by using the substitution method
3x-4y=29
y=x-8

Answer by mananth(16946)

3 x -5 y = 29 .............1
1 y = -8 + 1 x ---------2
/ 1 =
y = -8 + 1 x
Substitute y in (1)
3 x -5 ( -8 + 1 x ) = 29
3 x + 40 + -5 x = 29
3 x -5 x = 29 -40
-2 x = -11
/ -2
x= 5.5
Plug the value of x in (1)
3 x -5 y = 29
3 * 5.5 -5 y = 29
16.5 -5 y = 29
-5 y = 12.5
/ -5
y= -2.5
m.ananth@hotmail.ca